OpenStudy (anonymous):

square root of 12 open parentheses square root of 3 end root plus 2 close parentheses plus square root of 2 open parentheses square root of 8 end root minus square root of 6

OpenStudy (mertsj):

\[\sqrt{12}(\sqrt{3}+2)+\sqrt{2}(\sqrt{8}-\sqrt{6})\]

OpenStudy (mertsj):

Simplify sqrt12 and sqrt8 first.

OpenStudy (mertsj):

\[2\sqrt{3}(\sqrt{3}+2)+\sqrt{2}(2\sqrt{2}-\sqrt{6})\]

OpenStudy (mertsj):

Is that what you got?

OpenStudy (mertsj):

Now use the distributive property.

OpenStudy (anonymous):

is it 8 - 2 squart root of 3?

OpenStudy (mertsj):

\[2\sqrt{3}(\sqrt{3}+2)=6+4\sqrt{3}\] \[\sqrt{2}(2\sqrt{2}-\sqrt{6})=4-\sqrt{12}=4-2\sqrt{3}\] \[6+4\sqrt{3}+4-2\sqrt{3}=10+2\sqrt{3}\]

OpenStudy (anonymous):

Im confused what the next step should be

OpenStudy (mertsj):

There is no next step. 10+2sqrt3 is the final answer.

OpenStudy (anonymous):

oh i read it wrong sorry

OpenStudy (mertsj):

Not a problem.